A string is stretched between two fixed points at a distance 2l apart and the points of the string are given initial velocities v={CX/l, 0<x<l
C(2l-x)/l, l<x<l
Find the displacement of the string at any timr
1
Expert's answer
2020-08-20T17:46:14-0400
"Solution"


The displacement y(x,t) is given by the equation


"\frac{\delta^2y}{ \delta t^2}=a^2 \frac{\delta^2y}{ \delta x^2}------------(1)"

The suitable solution of (1) is given by

"y(x,t) = (A cos lx + B sin lx)(C cos lat + D sin lat) -------(2)"


The boundary conditions are

"(i)\ y(0,t) = 0,\ for\ t\ge0. \\
(ii)\ y(ℓ,t) ,\ for\ t\ge0. \\
(iii)\ y(x,0)=0\ for\ all\ 0<x<i\\
(iv)\ [\frac{\delta y}{\delta t}]_{t=0}=g(x)= \begin{cases}
 \frac{Cx}{l} & 0<x<l \\
 \frac{C(2l-x)}{l} & l<x<2l
 \end{cases}"

Using    (i) and (ii) in (2) , we get


"0 = A(Ccoslat + Dsinlat)" , for all "t \ge 0."

Therefore,           "A = 0"


Hence equation (2) becomes

"y(x,t) = B sinlx(Ccoslat + Dsinlat) ------------ (3)"

Using (ii) in (3), we get

 

"0 = Bsinl\u2113 (Ccoslat+Dsinlat)" , for all  "t \ge 0" , which gives "lℓ = n\pi"

Hence, "l=\frac{n \pi}{ℓ}" , n being an integer.


Thus

"y(x,t)=Bsin \frac{n \pi x}{l}[Ccos \frac{n \pi at}{l}+Dsin \frac{n \pi at}{l}]-----------(4)"


Using (iii) in (4), we get

"0=Bsin \frac{n \pi x}{l}C"


Therefore, "C = 0"

Hence "y(x,t)=Bsin \frac{n \pi x}{l} \cdot Dsin \frac{n \pi at}{l}=B_1sin \frac{n \pi x}{l} \cdot Dsin \frac{n \pi at}{l}"

Where "B_1=BD"


The most general solution is


"y(x,t)=\sum_{n=1}^\infin B_nsin \frac{n \pi x}{l} \cdot sin \frac{n \pi at}{l}---------(5)"


Differentiating (5) partially w.r.t t, we get


"\frac{\delta y}{\delta t}=\sum_{n=1}^\infin B_nsin \frac{n \pi x}{l} cos \frac{n \pi at}{l} \cdot \frac{n \pi a}{l}"


Using condition (iv) in the above equation, we get

"y_t(x,0)=g(x)= \begin{cases}
 \frac{Cx}{l} & 0<x<l \\
 \frac{2l-x}{l} & l<x<2l
\end{cases}"


Taking "\frac{Cx}{l},\ 0<x<l" we get


"\frac{Cx}{l}=\sum_{n=1}^\infin B_n \frac{n \pi a}{l} \cdot sin\frac{n \pi x}{l}\\ 
i.e\ B_n \frac{n \pi a}{l}=\frac{2}{l}\int_0^1 f(x)\cdot sin\frac{n \pi x}{l}dx\\
i.e\ B_n =\frac{2}{n \pi a}\int_0^1 f(x)\cdot sin\frac{n \pi x}{l}dx=\frac{2}{n \pi a}\int_0^1 \frac{Cx}{l}\cdot sin\frac{n \pi x}{l}dx= \frac{2}{n \pi a}\int_0^1 \frac{Cx}{l}d[\frac{-cos\frac{n \pi x}{l}}{\frac{n \pi x}{l}}]\\
= \frac{2}{n \pi a}\{\frac{Cx}{l}d[\frac{-cos\frac{n \pi x}{l}}{\frac{n \pi x}{l}}]-\frac{C}{l}[\frac{-sin\frac{n \pi}{l}}{\frac{n^2 \pi^2 }{l}}]\}=\frac{2}{n \pi a}\{\frac{-2cos\ n\pi+2}{\frac{n^3 \pi^3 }{l^3}}\}=\frac{2}{n \pi a} \cdot \frac{2l^3}{n^3 \pi^3}\{1-cos\ n\pi\}\\
i.e\ B_n =\frac{4l^3}{n^4 \pi^4 a}\{1-(-1)^n\}\\
or\\
B_n= \begin{cases}
 \frac{8l^3}{l} & if\ n\ is\ odd \\
 0 & if\ n\ is\ even
\end{cases}\\

\therefore\ solution\ is\\
y(x,t)=\frac{8l^3}{n^4a}\sum_{n=1}^\infin \frac{1}{(2n-1)^4}\sin\ \frac{(2n-1)\pi a t}{l}\sin\ \frac{(2n-1)\pi x}{l},\ \forall\ 0<x<l--->Answer"


Taking "\frac{C(2l-x)}{l},\ l<x<2l" we get


"\frac{2Cl-Cx}{l}=\sum_{n=1}^\infin B_n \frac{n \pi a}{2l} \cdot sin\frac{n \pi x}{2l}\\ 
i.e\ B_n \frac{n \pi a}{l}=\frac{2}{l}\int_l^{2l} g(x)\cdot sin\frac{n \pi x}{2l}dx\\
i.e\ B_n =\frac{2}{n \pi a}\int_l^{2l} g(x)\cdot sin\frac{n \pi x}{2l}dx=\frac{2}{n \pi a}\int_l^{2l} \frac{2Cl-Cx}{l}\cdot sin\frac{n \pi x}{2l}dx= \frac{2}{n \pi a}\int_l^{2l} \frac{2Cl-Cx}{l}d[\frac{-cos\frac{n \pi x}{2l}}{\frac{n \pi x}{2l}}]\\
= \frac{2}{n \pi a}\{\frac{2Cl-Cx}{l}d[\frac{-cos\frac{n \pi x}{2l}}{\frac{n \pi x}{2l}}]-\frac{2Cl-C}{l}[\frac{-sin\frac{n \pi}{2l}}{\frac{n^2 \pi^2 }{2l}}]\}=\frac{2}{n \pi a}\{\frac{-2cos\ n\pi+2}{\frac{n^3 \pi^3 }{4l^3}}\}=\frac{2}{n \pi a} \cdot \frac{8l^3}{n^3 \pi^3}\{1-cos\ n\pi\}\\
i.e\ B_n =\frac{16l^3}{n^4 \pi^4 a}\{1-(-1)^n\}\\
or\\
B_n= \begin{cases}
 \frac{32l^3}{l} & if\ n\ is\ odd \\
 0 & if\ n\ is\ even
\end{cases}\\

\therefore\ solution\ is\\
y(x,t)=\frac{32l^3}{n^4a}\sum_{n=1}^\infin \frac{1}{(2n-1)^4}\sin\ \frac{(2n-1)\pi a t}{l}\sin\ \frac{(2n-1)\pi x}{l},\ \forall\ l<x<2l---->Answer"